OpenStudy (anonymous):

Please help with how to verify/prove this identity. Also please show all work!! cos(x - π/4) = √(2)/2 •(cos x + sin x)

OpenStudy (anonymous):

\[\cos(x-y)=\cos x\cos y+\sin x\sin y\] Set \(y=\dfrac{\pi}{4}\).